Q# 2. Suppose you have the following information about a closed economy: C = 50 + 0.80 Y I =150 G = 175 Where C is consumption, I is investment, G is Government expenditure and Y is income.

a.      Find out the equilibrium level of income.

b.     Suppose I increases to 210 what is the new equilibrium level of income?

c.      What level of G is needed to achieve a target income of 4850?





1
Expert's answer
2021-04-06T06:56:28-0400

A closed economy is a type of economy where the import and export of goods and services don't happen, which implies that the economy is self-sufficient and has no trading activity from outside economics. The sole purpose of an open economy is to meet all the domestic consumers' needs within the country's border.

At equilibrium

i) "Y=C+I+G"

"Y=50+0.8Y+150+175"

"Y-0.8Y=375"

"0.2Y=""435"

"Y=1,875"

ii) If Investments increases to 210 then the new equilibrium becomes;

"Y=50+0.8Y+210+175"

"Y-0.8Y=435"

"0.2Y=435"

"Y=2,175"

iii) What level of G is needed to achieve a target income of 4850?

"Y=4,850"

Thus; "4850=50+0.8(4850)+150+G"

"4850=200+3880+G"

"G=4850-4080"

"G=770"